OpenStudy (scarlettfarra2000):

32/20= 24/? (proportion)

OpenStudy (anonymous):

\[\frac{ 32 }{ 20 }=\frac{ 24 }{ x }\]

OpenStudy (anonymous):

multiplyboth sides by 20 to clear the 20 from the left side

OpenStudy (anonymous):

\[32=(\frac{ 24 }{ x })(\frac{ 20 }{1 })\]

OpenStudy (texaschic101):

32/20 = 24/x to solve proportions, you must first cross multiply, then solve for the variable. (32)(x) = (20)(24) now, just do the multiplying...and solve for x. A little tip...proportions are nothing but equivalent fractions.

OpenStudy (anonymous):

\[32=\frac{ 480 }{ x }\]

OpenStudy (anonymous):

32x=480

OpenStudy (anonymous):

\[x=\frac{ 480 }{ 32 }\]

OpenStudy (anonymous):

i leave you with the easiest part

OpenStudy (texaschic101):

just divide 480 by 32 and you will have x

OpenStudy (scarlettfarra2000):

x=15

OpenStudy (texaschic101):

yep...x = 15